# RepLoRA: Reparameterizing Low-Rank Adaptation via the Perspective of Mixture of Experts

### v1

@@ 4.2 With Reparametrization: Optimal Sample Complexity
where $\varphi_{1}:\mathbb{R}^{m\times m}\to\mathbb{R}^{r\times\bar{d}}$, $\varphi_{2}:\mathbb{R}^{m\times m}\to\mathbb{R}^{r\times\bar{d}}$ are some functions, ${\bm{A}}\in\mathbb{R}^{m\times m},{\bm{B}}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$ are learnable matrices with given dimensions $m,m^{\prime}\geq 1$. We specifically note for the simplicity of the theoretical development, this formulation is simplified compared to what was used in practice because we set the low-rank matrices of queries to be equal to that of values. Nevertheless, as we will show in this section, even with this simplified formulation, reparameterization gives superior sample complexity compared to vanilla LoRA without reparameterization. After training, the reparameterization can be discarded, and we only need to store the low-rank matrices ${\bm{A}}_{Q},{\bm{A}}_{V},{\bm{B}}_{Q},{\bm{B}}_{V}$. We observe that this reparameterization technique encodes a shared structure between the query and value low-rank matrices. The primary difference compared to the original LoRA is that instead of learning the low-rank matrices separately, we reparameterize those matrices as the output of the two shared structures $\varphi_{1},\varphi_{2}$. The subsequent sections will present the theoretical advantage of reparameterization based on the following two settings:
(i) Linear reparametrization: $\varphi_{1}({{\bm{A}}})=\bm{W}_{1}{\bm{A}}$ and $\varphi_{2}({{\bm{B}}})=\bm{W}_{2}{\bm{B}}$.
@@ 4.2.1 Simple Linear Reparameterization
$\sum_{j=1}^{L}\frac{\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\bm{W}_{2,j}^{*}{\bm{B}}_{j}^{*}\bm{W}_{1,j}^{*}{\bm{A}}_{j}^{*}){\bm{M}}^{0}_{K}\mathbb{X}+c^{*}_{j})}{\bar{D}_{f}(\mathbb{X})}\cdot({\bm{M}}^{0}_{V}+\bm{W}_{2,j}^{*}{\bm{B}}_{j}^{*}\bm{W}_{1,j}^{*}{\bm{A}}_{j}^{*})\mathbb{X},$ (13)
where we have $\bar{D}_{f}(\mathbb{X})=\sum_{k=1}^{L}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\bm{W}_{2,k}^{*}{\bm{B}}_{k}^{*}\bm{W}_{1,k}^{*}{\bm{A}}_{k}^{*}){\bm{M}}^{0}_{K}\mathbb{X}+c^{*}_{k})$, while the mixing measure is of the form $\bar{G}_{*}:=\sum_{j^{\prime}=1}^{L}\exp(c^{*}_{j^{\prime}})\delta_{\bm{W}_{2,j^{\prime}}^{*}{\bm{B}}_{j^{\prime}}^{*}\bm{W}_{1,j^{\prime}}^{*}{\bm{A}}_{j^{\prime}}^{*}}$. Similar to Section 4.1, we estimate the low-rank matrices via estimating the ground-truth mixing measure $\bar{G}_{*}$ using the least square method:
$\bar{G}_{n}:=\textnormal{arg}\min_{\bar{G}\in\bar{\mathcal{G}}_{L^{\prime}}(\Theta)}\sum_{i=1}^{n}\Big{\|}{\bm{Y}}_{i}-f_{\bar{G}}(\mathbb{X}_{i})\Big{\|}^{2},$ (14)
@@
$+{\sum_{j^{\prime}\in[L]:|\mathcal{V}_{j^{\prime}}|=1}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(c_{i})\|{\bm{Z}}_{i}-{\bm{Z}}_{j^{\prime}}^{*}\|}+\sum_{j^{\prime}\in[L]:|\mathcal{V}_{j^{\prime}}|>1}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(c_{i})\|{\bm{Z}}_{i}-{\bm{Z}}_{j^{\prime}}^{*}\|^{2},$
@@ 4.2.2 Non-Linear Reparameterization
$\sum_{j=1}^{N}\frac{\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}(\bm{W}^{*}_{2,j}{\bm{B}}^{*}_{j})\sigma_{1}(\bm{W}^{*}_{1,j}{\bm{A}}^{*}_{j})){\bm{M}}^{0}_{K}\mathbb{X}+c^{*}_{j})}{\widetilde{D}_{f}(\mathbb{X})}\cdot({\bm{M}}^{0}_{V,j}+\sigma_{2}(\bm{W}^{*}_{2,j}{\bm{B}}^{*}_{j})\sigma_{1}(\bm{W}^{*}_{1,j}{\bm{A}}^{*}_{j}))\mathbb{X},$ (15)
@@
$\widetilde{G}_{n}:=\textnormal{arg}\min_{\widetilde{G}\in\widetilde{\mathcal{G}}_{L^{\prime}}(\Theta)}\sum_{i=1}^{n}\Big{\|}{\bm{Y}}_{i}-f_{\widetilde{G}}(\mathbb{X}_{i})\Big{\|}^{2}.$ (16)
@@
$\mathcal{D}_{3}(\widetilde{G},\widetilde{G}_{*}):=\sum_{j^{\prime}=1}^{L}\Big{|}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(c_{i})-\exp(c^{*}_{j^{\prime}})\Big{|}+\sum_{\begin{subarray}{c}j^{\prime}\in[L]:|\mathcal{V}_{j^{\prime}}|=1,i\in\mathcal{V}_{j^{\prime}}\end{subarray}}\exp(c_{i})(\|\Delta(\bm{W}_{2}{\bm{B}})_{ij^{\prime}}\|+\|\Delta(\bm{W}_{1}{\bm{A}})_{ij^{\prime}}\|)$
$\hskip 50.00008pt+\sum_{\begin{subarray}{c}j^{\prime}\in[L]:|\mathcal{V}_{j^{\prime}}|>1,i\in\mathcal{V}_{j^{\prime}}\end{subarray}}\exp(c_{i})(\|\Delta(\bm{W}_{2}{\bm{B}})_{ij^{\prime}}\|^{2}+\|\Delta(\bm{W}_{1}{\bm{A}})_{ij^{\prime}}\|^{2}),$
where we denote $\Delta(\bm{W}_{2}{\bm{B}})_{ij^{\prime}}:=\bm{W}_{2,i}{\bm{B}}_{i}-\bm{W}^{*}_{2,j^{\prime}}{\bm{B}}^{*}_{j^{\prime}}$ and $\Delta(\bm{W}_{1}{\bm{A}})_{ij^{\prime}}:=\bm{W}_{1,i}{\bm{A}}_{i}-\bm{W}^{*}_{1,j^{\prime}}{\bm{A}}^{*}_{j^{\prime}}$ for any $i,j^{\prime}$. Before presenting the main result of this section in Theorem 4.3, it is necessary to impose some mild assumptions on the activations $\sigma_{1}$ and $\sigma_{2}$.
Assumptions. We have the following assumptions on the activation functions $\sigma_{1}$ and $\sigma_{2}$:
(A.2) (Uniform Lipschitz) Let $F(\mathbb{X};{\bm{B}},{\bm{A}}):=\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}})\sigma_{1}({\bm{A}}))\mathbb{X})({\bm{M}}^{0}_{V}+\sigma_{2}({\bm{B}})\sigma_{1}({\bm{A}}))\mathbb{X}$. Then, for any coefficient $\tau\in\{1,2\}$, the following inequality holds
$\sum_{|\alpha|=\tau}\Bigg{|}\Big{(}\frac{\partial^{|\alpha|}F}{\partial{\bm{A}}^{\alpha_{1}}\partial{\bm{B}}^{\alpha_{2}}}(\mathbb{X};{\bm{B}},{\bm{A}})$ $-\frac{\partial^{|\alpha|}F}{\partial{\bm{A}}^{\alpha_{1}}\partial{\bm{B}}^{\alpha_{2}}}(\mathbb{X};{\bm{B}}^{\prime},{\bm{A}}^{\prime})\Big{)}\gamma^{\alpha}\Bigg{|}\leq C\|({\bm{B}},{\bm{A}})-({\bm{B}}^{\prime},{\bm{A}}^{\prime})\|^{\zeta}\|\gamma\|^{\tau},$
for any vector $\gamma\in\mathbb{R}^{2dr}$ and for some positive constants $\zeta$ and $C$ that are independent of $\mathbb{X}$ and $({\bm{B}},{\bm{A}}),({\bm{B}}^{\prime},{\bm{A}}^{\prime})$. Finally, in the summation, the notation $\alpha=(\alpha_{1},\alpha_{2})\in\mathbb{N}^{r\times d}\times\mathbb{N}^{d\times r}$.
are linearly independent for almost surely ${\bm{X}}$.
@@
Assume that the activation functions $\sigma_{1}$ and $\sigma_{2}$ satisfy the Assumptions (A.1)-(A.3). Then, the least square estimator $\widetilde{G}_{n}$ converges to the true mixing measure $\widetilde{G}_{*}$ at the following rate:
@@ 5 Reparameterized Low-Rank Adaptation
where ${\bm{A}},{\bm{B}}$ are learnable matrices, and $g_{\theta_{\bm{A}}},g_{\theta_{\bm{B}}}$ are two-layer MLPs with a shared part and distinct output heads. In this approach, ${\bm{A}}_{Q}$ and ${\bm{A}}_{V}$ are derived from a shared underlying input ${\bm{A}}$, with distinct outputs ${\bm{A}}_{Q}$ and ${\bm{A}}_{V}$ produced by the separated heads of $g_{{\bm{\theta}}_{\bm{A}}}$. Similarly, ${\bm{B}}_{Q}$ and ${\bm{B}}_{V}$ follow the same structure, leveraging a shared ${\bm{B}}$ input. While we focus on fine-tuning the queries and values to streamline the analysis, this formulation can naturally be extended to fine-tune the keys as well. We implement ${\bm{A}}$ and ${\bm{B}}$ as diagonal matrices to ensure parameter efficiency. After training, the reparameterization $g_{{\bm{\theta}}_{\bm{A}}}$ and $g_{{\bm{\theta}}_{\bm{B}}}$ can be discarded, and only the fine-tuned matrices ${\bm{A}}_{Q}$, ${\bm{A}}_{V}$, ${\bm{B}}_{Q}$, and ${\bm{B}}_{V}$ need to be retained for inference. Hence, this approach does not incur any additional computational overhead for inference. An illustration of this method is provided in Figure 1.
@@ 6 Experiments
Settings. We extensively experiment across multiple domains to demonstrate the versatility and effectiveness of RepLoRA on a wide range of tasks. Our evaluation spans four distinct settings: language (commonsense reasoning), image (classification), video (video action recognition), and multi-modal (image/video-text understanding). To provide a comprehensive evaluation, we compare RepLoRA against several PEFT methods, such as Full Fine-tuning, Prefix Tuning [29], LoRA [12], and Series Adapter [11]. As summarized in Figure 3, RepLoRA consistently outperforms LoRA across all settings, highlighting its robust adaptability and superior performance across diverse tasks.
Metrics. In Parameter-Efficient Fine-Tuning (PEFT), evaluations typically focus on performance and the number of trainable parameters. The goal is to maximize performance while minimizing the parameters required. To assess this trade-off, in addition to reporting performance, we adopt the Performance-Parameter Trade-off (PPT) metric, proposed by [53]. Specifically, for a PETL algorithm $M$, the PPT metric incorporates its task performance $M_{t}$, the number of trainable parameters $P_{M}$, and a normalization constant $C$. Formally, we have:
@@
Commonsense Reasoning. In our first experiment, we compare the performance of RepLoRA against LoRA and other PEFT methods using LLaMA-7B/13B [48] on the Commonsense Reasoning task. We also include the accuracy of ChatGPT, measured with the GPT-3.5-turbo API with a zero-shot Chain of Thought approach [52]. The commonsense reasoning benchmark consists of eight sub-tasks with predefined training and testing datasets. Following the settings outlined in [13], we combine the training datasets from all eight sub-tasks into a single training dataset and evaluate performance on the individual testing datasets for each task. To ensure a fair comparison, we fine-tuned the models with RepLoRA using the same configuration as LoRA, keeping the rank fixed. As presented in Table 1, RepLoRA achieves significantly better results than LoRA across all settings, delivering substantial improvements not only in accuracy but also in the PPT score, emphasizing its parameter efficiency.
Image Classification. We extend our evaluation of RepLoRA to the image domain and fine-tune the ViT-B/16 architecture [3], pre-trained on the ImageNet-21K dataset [2], on two challenging benchmarks: the VTAB-1K dataset suite [55] and the FGVC dataset collection [16].
@@
Video Action Recognition. Given RepLoRA’s strong performance in the image domain, we expand our experiments to the video domain. We evaluate our method against baseline approaches using the Video Swin Transformer on two datasets: SSv2 [7], which offers a rich dataset with abundant data, and HMDB51 [19], which presents a more challenging scenario with limited data and fewer categories. Despite the contrasting characteristics of these datasets, Table 4 indicates that RepLoRA remarkably outperforms all baselines while maintaining parameter efficiency, underscoring its adaptability and robustness across diverse data settings.
Image/Video-Text understanding. Having demonstrated that RepLoRA outperforms the baselines on the language and vision tasks, we attempt to see if RepLoRA remains competitive on multi-modality tasks. This experiment compares RepLoRA with LoRA and fully fine-tune (FT) on the VL-BART [25]. The experiments were conducted on four image-text tasks: $\texttt{VQA}^{\texttt{v}^{2}}$ [8], GQA [14] for vision question-answering, NLVR2 [45] for visual reasoning, MSCOCO [1] for image captioning, and four video-text tasks from the VALUE benchmark [28]: TVQA [22] and How2QA [27] for video question answering, TVC [23] and YC2C [57] for video captioning. We follow [46] and adopt the same setup of LoRA when applying RepLoRA. It is evident that RepLoRA consistently surpasses both FT and LoRA in accuracy and PPT in both Tables 6 and Table 6. In particular, RepLoRA exceeds LoRA’s performance by nearly 2% in image-text understanding tasks and roughly 4% in video-text understanding tasks, reaching the performance of FT.
Table 5: Image-text understanding performance on VL-BART.
@@
Table 6: Video-text understanding performance on VL-BART.
@@
Enhancing Sampling Efficiency. Our theoretical analysis has demonstrated that reparameterizing LoRA achieves superior rates of sample efficiency compared to vanilla LoRA. To empirically validate this claim, we evaluate the sample efficiency of RepLoRA on FGVC datasets. Following the approach of [5], we subsample each class at fractions $f=\{1\%,10\%,30\%,50\%,100\%\}$ and scale the number of training epochs by $1/f$, ensuring the total number of data seen by the model remains constant. Figure 2 shows that RepLoRA consistently outperforms LoRA across all sampling fractions. The improvements are significant at smaller fractions, with RepLoRA achieving a remarkable $40.4\%$ gap at $f=1\%$. Moreover, we emphasize that RepLoRA matches LoRA’s performance with only $30\%$ training fraction, therefore underscoring RepLoRA’s superior sample efficiency, as predicted by our theoretical analysis. We refer to Appendix D.1 for a breakdown of these results.
Linear vs. Non-linear Reparameterization. Another conclusion from the theoretical analysis is that even a simple linear reparameterization with a shared structure offers significant efficiency gains compared to vanilla LoRA. Furthermore, as shown in Theorems 4.2 and 4.3, incorporating non-linear reparameterization further improves the rate of low-rank matrix estimation. To validate this hypothesis, we conducted empirical experiments, with the results presented in Figure 3. These results demonstrate that non-linear reparameterization outperforms the linear setting by substantial margins, underscoring its effectiveness. For a more detailed comparison of linear versus non-linear reparameterization, please refer to Appendix D.2.

### v2

@@ 5.2 With Reparametrization: Optimal Sample Complexity
where $\varphi_{1}:\mathbb{R}^{m\times m}\to\mathbb{R}^{r\times\bar{d}}$, $\varphi_{2}:\mathbb{R}^{m\times m}\to\mathbb{R}^{r\times\bar{d}}$ are some functions, ${\bm{A}}\in\mathbb{R}^{m\times m},{\bm{B}}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$ are learnable matrices with given dimensions $m,m^{\prime}\geq 1$. We specifically note that, for the simplicity of the theoretical development, this formulation is simplified compared to what was used in practice because we set the low-rank matrices of queries to be equal to that of values. As we will show in this section, even with this simplified formulation, reparameterization gives superior sample complexity compared to vanilla LoRA without reparameterization.
After training, the reparameterization can be discarded, and only the low-rank matrices ${\bm{A}}_{Q},{\bm{A}}_{V},{\bm{B}}_{Q},{\bm{B}}_{V}$ need to be stored. We observe that the reparameterization strategy implicitly encodes a shared structure between the query and value low-rank matrices. The primary difference compared to the original LoRA is that instead of learning the low-rank matrices separately, we reparameterize those matrices as the output of the two shared structures $\varphi_{1},\varphi_{2}$. To study the theoretical advantages of the reparametrization technique, we focus on the following two settings of the functions $\varphi_{1}$ and $\varphi_{2}$:
(i) Simple linear reparametrization: $\varphi_{1}({{\bm{A}}})=\bm{W}_{1}{\bm{A}}$ and $\varphi_{2}({{\bm{B}}})=\bm{W}_{2}{\bm{B}}$.
@@
It should be noted that the above reparametrization settings can totally be generalized to the scenarios where the matrices $A_{Q}$ and $A_{V}$ share only the learnable matrix $A$. In particular, we can reformulate those matrices as $A_{Q}=W_{Q,1}A$ and $A_{V}=W_{V,1}A$ for the simple linear reparametrization and as $A_{Q}=\sigma_{1}(W_{Q,1}A)$ and $A_{V}=\sigma_{1}(W_{V,1}A)$ for the non-linear reparametrization. In order to tailor to these settings, it is necessary to include several terms involving parameters $W_{Q,1}$ and $W_{V,1}$ rather than merely parameters $W_{1}$ as in the above settings. However, we realize that these terms not only provide no additional information on our convergence analysis but also make it unnecessarily complicated. Therefore, we assume without loss of generalization that $A_{Q}=A_{V}=W_{1}A$ or $A_{Q}=A_{V}=\sigma_{1}(W_{1}A)$ to simplify the analysis, making it more accessible.
@@ 5.2.1 Simple Linear Reparameterization
$\sum_{j=1}^{L}\frac{\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\bm{W}_{2,j}^{*}{\bm{B}}_{j}^{*}\bm{W}_{1,j}^{*}{\bm{A}}_{j}^{*}){\bm{M}}^{0}_{K}\mathbb{X}+c^{*}_{j})}{\bar{D}_{f}(\mathbb{X})}$
$\hskip 71.13188pt\cdot({\bm{M}}^{0}_{V}+\bm{W}_{2,j}^{*}{\bm{B}}_{j}^{*}\bm{W}_{1,j}^{*}{\bm{A}}_{j}^{*})\mathbb{X},$ (13)
where we denote $\bar{D}_{f}(\mathbb{X})=\sum_{k=1}^{L}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\bm{W}_{2,k}^{*}{\bm{B}}_{k}^{*}\bm{W}_{1,k}^{*}{\bm{A}}_{k}^{*}){\bm{M}}^{0}_{K}\mathbb{X}+c^{*}_{k})$, while the mixing measure is of the form $\bar{G}_{*}:=\sum_{j^{\prime}=1}^{L}\exp(c^{*}_{j^{\prime}})\delta_{\bm{W}_{2,j^{\prime}}^{*}{\bm{B}}_{j^{\prime}}^{*}\bm{W}_{1,j^{\prime}}^{*}{\bm{A}}_{j^{\prime}}^{*}}$. Similar to Section 5.1, we estimate the unknown low-rank matrices via estimating the ground-truth mixing measure $\bar{G}_{*}$ using the least square method:
$\bar{G}_{n}\in\textnormal{arg}\min_{\bar{G}\in\bar{\mathcal{G}}_{L^{\prime}}(\Theta)}\sum_{i=1}^{n}\Big{(}{\bm{Y}}_{i}-f_{\bar{G}}(\mathbb{X}_{i})\Big{)}^{2},$ (14)
@@
$+{\sum_{j^{\prime}\in[L]:|\mathcal{V}_{j^{\prime}}|=1}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(c_{i})\|{\bm{Z}}_{i}-{\bm{Z}}_{j^{\prime}}^{*}\|}$
$+\sum_{j^{\prime}\in[L]:|\mathcal{V}_{j^{\prime}}|>1}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(c_{i})\|{\bm{Z}}_{i}-{\bm{Z}}_{j^{\prime}}^{*}\|^{2},$
@@ 5.2.2 Non-Linear Reparameterization
$\sum_{j=1}^{N}\frac{\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}(\bm{W}^{*}_{2,j}{\bm{B}}^{*}_{j})\sigma_{1}(\bm{W}^{*}_{1,j}{\bm{A}}^{*}_{j})){\bm{M}}^{0}_{K}\mathbb{X}+c^{*}_{j})}{\widetilde{D}_{f}(\mathbb{X})}$
$\hskip 42.67912pt\cdot({\bm{M}}^{0}_{V,j}+\sigma_{2}(\bm{W}^{*}_{2,j}{\bm{B}}^{*}_{j})\sigma_{1}(\bm{W}^{*}_{1,j}{\bm{A}}^{*}_{j}))\mathbb{X},$ (15)
@@
$\widetilde{G}_{n}\in\textnormal{arg}\min_{\widetilde{G}\in\widetilde{\mathcal{G}}_{L^{\prime}}(\Theta)}\sum_{i=1}^{n}\Big{(}{\bm{Y}}_{i}-f_{\widetilde{G}}(\mathbb{X}_{i})\Big{)}^{2}.$ (16)
@@
$\mathcal{D}_{3}(\widetilde{G},\widetilde{G}_{*}):=\sum_{j^{\prime}=1}^{L}\Big{|}\sum_{i\in\mathcal{V}_{j^{\prime}}}\exp(c_{i})-\exp(c^{*}_{j^{\prime}})\Big{|}$
$+\sum_{\begin{subarray}{c}j^{\prime}\in[L]:\\ |\mathcal{V}_{j^{\prime}}|=1,\\ i\in\mathcal{V}_{j^{\prime}}\end{subarray}}\exp(c_{i})(\|\Delta(\bm{W}_{2}{\bm{B}})_{ij^{\prime}}\|+\|\Delta(\bm{W}_{1}{\bm{A}})_{ij^{\prime}}\|)$
$+\sum_{\begin{subarray}{c}j^{\prime}\in[L]:\\ |\mathcal{V}_{j^{\prime}}|>1,\\ i\in\mathcal{V}_{j^{\prime}}\end{subarray}}\exp(c_{i})(\|\Delta(\bm{W}_{2}{\bm{B}})_{ij^{\prime}}\|^{2}+\|\Delta(\bm{W}_{1}{\bm{A}})_{ij^{\prime}}\|^{2}),$
where we denote $\Delta(\bm{W}_{2}{\bm{B}})_{ij^{\prime}}:=\bm{W}_{2,i}{\bm{B}}_{i}-\bm{W}^{*}_{2,j^{\prime}}{\bm{B}}^{*}_{j^{\prime}}$ and $\Delta(\bm{W}_{1}{\bm{A}})_{ij^{\prime}}:=\bm{W}_{1,i}{\bm{A}}_{i}-\bm{W}^{*}_{1,j^{\prime}}{\bm{A}}^{*}_{j^{\prime}}$ for any $i,j^{\prime}$. Before presenting the main result of this section in Theorem 5.3, it is necessary to impose some mild assumptions on the activations $\sigma_{1}$ and $\sigma_{2}$. Due to the space limit, we defer those assumptions to the proof of Theorem 5.3 in Appendix A.3.
@@
Assume that the activation functions $\sigma_{1}$ and $\sigma_{2}$ meet the assumptions specified in Appendix A.3. Then, the estimator $\widetilde{G}_{n}$ converges to the true mixing measure $\widetilde{G}_{*}$ at the following rate:
@@ 6 Reparameterized Low-rank Adaptation
where ${\bm{A}},{\bm{B}}$ are learnable matrices, and $g_{\theta_{\bm{A}}},g_{\theta_{\bm{B}}}$ are two-layer MLPs with a shared part and distinct output heads. In this approach, ${\bm{A}}_{Q}$ and ${\bm{A}}_{V}$ are derived from a shared underlying input ${\bm{A}}$, with distinct outputs ${\bm{A}}_{Q}$ and ${\bm{A}}_{V}$ produced by the separated heads of $g_{{\bm{\theta}}_{\bm{A}}}$. Similarly, ${\bm{B}}_{Q}$ and ${\bm{B}}_{V}$ follow the same structure, leveraging a shared ${\bm{B}}$ input. While we focus on fine-tuning the queries and values to streamline the analysis, this formulation can naturally be extended to fine-tune the keys. We implement ${\bm{A}}$ and ${\bm{B}}$ as diagonal matrices to ensure parameter efficiency. After training, the reparameterization $g_{{\bm{\theta}}_{\bm{A}}}$ and $g_{{\bm{\theta}}_{\bm{B}}}$ can be discarded, and only the fine-tuned matrices ${\bm{A}}_{Q}$, ${\bm{A}}_{V}$, ${\bm{B}}_{Q}$, and ${\bm{B}}_{V}$ need to be retained for inference. Hence, this approach does not incur any additional computational overhead for inference. An illustration of this method is provided in Figure 1.
@@ 7 Experiments
Experimental Settings. We conduct extensive experiments across multiple domains to demonstrate the versatility and effectiveness of RepLoRA in a wide range of tasks. Our evaluation spans four distinct settings: language (commonsense reasoning), image (classification), video (video action recognition), and multi-modal (image/video-text understanding). To provide a comprehensive evaluation, we compare RepLoRA against several PEFT methods, such as Full Fine-tuning, Prefix Tuning (Li & Liang, 2021), LoRA (Hu et al., 2022), DoRA (Liu et al., 2024b), and Series Adapter (Houlsby et al., 2019). As summarized in Figure 3, RepLoRA consistently outperforms LoRA across all settings, highlighting its robust adaptability and superior performance in various tasks.
Evaluation Metrics. In Parameter-Efficient Fine-Tuning (PEFT), evaluations typically focus on performance and the number of trainable parameters. The goal is to maximize performance while minimizing the parameters required. To assess this trade-off, in addition to reporting performance, we adopt the Performance-Parameter Trade-off (PPT) metric, proposed by Xin et al. (2024). Specifically, for a PETL algorithm $M$, the PPT metric incorporates its task performance $M_{t}$, the number of trainable parameters $P_{M}$, and a normalization constant $C$. Formally, we have:
@@
Commonsense Reasoning. In our first experiment, we compare the performance of RepLoRA against LoRA and other PEFT methods using LLaMA-7B/13B (Touvron et al., 2023) on the Commonsense Reasoning task. We also include the accuracy of ChatGPT, measured with the GPT-3.5-turbo API with a zero-shot Chain of Thought approach (Wei et al., 2022). The commonsense reasoning benchmark comprises eight sub-tasks with predefined training and testing datasets. Following the settings outlined in (Hu et al., 2023), we combine the training datasets from all eight sub-tasks into a single training dataset and evaluate performance on the individual testing datasets for each task. To ensure a fair comparison, we fine-tuned the models with RepLoRA using the same configuration as LoRA, keeping the rank fixed. As presented in Table 1, RepLoRA achieves significantly better results than LoRA across all settings, delivering substantial improvements not only in accuracy but also in the PPT score, emphasizing its parameter efficiency.
Image Classification. We extend our evaluation of RepLoRA to the image domain and fine-tune the ViT-B/16 architecture (Dosovitskiy, 2021), pre-trained on the ImageNet-21K dataset (Deng et al., 2009), on two challenging benchmarks: the VTAB-1K dataset suite (Zhai et al., 2019) and the FGVC dataset collection (Jia et al., 2022).
@@
Video Action Recognition. Given RepLoRA’s strong performance in the image domain, we expand our experiments to the video domain. We evaluate our method against baseline approaches using the Video Swin Transformer on two datasets: SSv2 (Goyal et al., 2017a), which offers a rich dataset with abundant data, and HMDB51 (Kuehne et al., 2011), which presents a more challenging scenario with limited data and fewer categories. Despite the contrasting characteristics of these datasets, Table 4 indicates that RepLoRA remarkably outperforms all baselines while maintaining parameter efficiency, underscoring its adaptability and robustness across diverse data settings.
@@
Image/Video-Text understanding. Having demonstrated that RepLoRA outperforms the baselines on the language and vision tasks, we attempt to see if RepLoRA remains competitive on multi-modality tasks. This experiment compares RepLoRA with LoRA and full fine-tuning (FT) on the VL-BART (Lewis et al., 2020). The experiments were conducted on four image-text tasks: $\texttt{VQA}^{\texttt{v}^{2}}$ (Goyal et al., 2017b), GQA (Hudson & Manning, 2019) for vision question-answering, NLVR2 (Suhr et al., 2019) for visual reasoning, MSCOCO (Chen et al., 2015) for image captioning, and four video-text tasks from the VALUE benchmark (Li et al., 2021): TVQA (Lei et al., 2018) and How2QA (Li et al., 2020) for video question answering, TVC (Lei et al., 2020) and YC2C (Zhou et al., 2018) for video captioning. We follow Sung et al. (2022) and adopt the same setup of LoRA when applying RepLoRA. It is evident that RepLoRA consistently surpasses both FT and LoRA in accuracy and PPT in both Tables 5 and Table 6. In particular, RepLoRA exceeds LoRA’s performance by nearly 2% in image-text understanding tasks and roughly 4% in video-text understanding tasks, reaching the performance of FT.
Table 5: Performance on image-text tasks with VL-BART.
@@
Table 6: Performance on video-text tasks with VL-BART.
@@
Enhancing Sampling Efficiency. Our theoretical analysis has demonstrated that reparameterization achieves superior rates of sample efficiency compared to vanilla LoRA. To validate this claim, we evaluate the sample efficiency of RepLoRA on FGVC datasets. Following the approach of d’Ascoli et al. (2021), we subsample each class at fractions $f=\{1\%,10\%,30\%,50\%,100\%\}$ and scale the number of training epochs by $1/f$, ensuring the total number of data seen by the model remains constant. Figure 2 shows that RepLoRA consistently outperforms LoRA across all sampling fractions. The improvements are significant at smaller fractions, with RepLoRA achieving a remarkable 40.4% gap at $f=1\%$. More importantly, we emphasize that RepLoRA matches LoRA’s performance with only 30% training fraction, therefore underscoring RepLoRA’s superior sample efficiency, as predicted by our theoretical analysis. We refer to Appendix C.1 for a breakdown of these results.
Linear vs. Non-linear Reparameterization. Another conclusion from the theoretical analysis is that even a simple linear reparameterization with a shared structure offers significant efficiency gains compared to vanilla LoRA. Furthermore, as shown in Theorems 5.2 and 5.3, incorporating non-linear reparameterization further improves the rate of low-rank matrix estimation. To validate this hypothesis, we conducted empirical experiments, with the results presented in Figure 3. These results demonstrate that non-linear reparameterization outperforms the linear setting by substantial margins, underscoring its effectiveness. For a more detailed comparison of linear versus non-linear reparameterization performance, please refer to Appendix C.2.